### v1

# Leveraging Large Language Models to Measure Gender Bias in Gendered Languages

Abstract
Gender bias in text corpora used in various natural language processing (NLP) contexts, such as for training large language models (LLMs), can lead to the perpetuation and amplification of societal inequalities. This is particularly pronounced in gendered languages like Spanish or French, where grammatical structures inherently encode gender, making the bias analysis more challenging. Existing methods designed for English are inadequate for this task due to the intrinsic linguistic differences between English and gendered languages. This paper introduces a novel methodology that leverages the contextual understanding capabilities of LLMs to quantitatively analyze gender representation in Spanish corpora. By utilizing LLMs to identify and classify gendered nouns and pronouns in relation to their reference to human entities, our approach provides a nuanced analysis of gender biases. We empirically validate our method on four widely-used benchmark datasets, uncovering significant gender disparities with a male-to-female ratio ranging from 4:1 to 6:1. These findings demonstrate the value of our methodology for bias quantification in gendered languages and suggest its application in NLP, contributing to the development of more equitable language technologies.
Leveraging Large Language Models to Measure
Gender Bias in Gendered Languages
Erik Derner1, Sara Sansalvador de la Fuente1, Yoan Gutiérrez2, Paloma Moreda2, Nuria Oliver1 1 ELLIS Alicante, Alicante, Spain  2 University of Alicante, Alicante, Spain Correspondence: erik@ellisalicante.org
1 Introduction
Figure 1: Overview of the proposed method for gender bias detection and measurement in gendered languages using LLMs.
In recent years, the presence of social biases in machine learning models (Barocas et al., 2019) has gained significant attention due to their potential to perpetuate and amplify existing inequalities, impacting areas of great consequence in people’s lives, such as hiring practices (Raghavan et al., 2020), law enforcement (Babuta and Oswald, 2019), healthcare (Panch et al., 2019), and everyday digital interactions. Among various forms of bias, gender bias is particularly concerning because it affects roughly half of the global population and has pervasive effects across different sectors of society. This concern is amplified in the area of natural language processing (NLP), where biases present in training corpora can propagate through machine learning models and impact a broad range of applications.
In this work, we consider bias in text corpora used to train large language models (LLMs). The main motivation for studying bias in training data of LLMs stems from the lifecycle of these models. From training to deployment, biases in the training data can not only be perpetuated but even amplified by the models (Kotek et al., 2023). This data often comes from diverse sources, including books, websites, social media, and other digital texts. If these sources contain biases – which they frequently do, reflecting societal prejudices and stereotypes – the models learn these biases as part of their training process. Hence, the first stage – data curation – becomes of paramount importance. It involves selecting, cleaning, and annotating the training data to minimize biases as much as possible. Effective data curation requires a nuanced understanding of the types of biases that can exist and the methods to mitigate them, including removing or balancing biased content, ensuring diverse and representative data samples, and incorporating fairness and ethical guidelines in the data selection process (Gallegos et al., 2024).
However, existing methods for evaluating bias in text, such as gender polarity (Dhamala et al., 2021) and co-occurrences (Bordia and Bowman, 2019), have been primarily developed for English. These approaches often rely on the neutrality of the English language in terms of grammatical gender, allowing for relatively straightforward identification and categorization of gendered expressions. Nonetheless, such an approach is not directly applicable to gendered languages like Spanish, where nouns, articles, adjectives and other word forms reflect gender, making the detection and analysis of gender bias a much harder problem to solve.
Given the limitations of existing methods and the prevalence of gendered languages in the world, the focus of this paper is to develop a novel and robust method to detect gender bias in corpora used to train LLMs in Spanish, a representative gendered language spoken by over 500 million people across diverse cultures and regions in the world. In Spanish, every noun carries a gender marker but only the words that refer to human beings are reflective of a potential gender bias.
We propose a novel methodology that leverages state-of-the-art LLMs to detect and analyze gender bias in Spanish text corpora. Unlike traditional methods that rely on part-of-speech tagging or specific machine learning models to detect person references, our approach benefits from the contextual understanding capabilities of LLMs. By prompting LLMs to identify and classify nouns and pronouns by their reference to persons and their grammatical gender, we introduce a method that respects the linguistic structure of Spanish while providing the necessary depth of analysis.
We first provide an overview of the most relevant literature on biases in NLP systems in English and non-English languages, followed by a detailed description of the proposed methodology. We empirically evaluate the proposed method on several benchmark datasets and uncover significant levels of gender imbalance with a male:female ratio ranging between 4:1 and 6:1, depending on the dataset. Interestingly, the English translations of the same corpora exhibit a male:female gender ratio of 1:1 to 3.5:1. Our results illustrate the importance of developing novel methods to detect gender bias in text corpora in gendered languages, such as Spanish. We conclude with a discussion, limitations, ethical implications, and lines of future work.
2 Related Work
Biases in NLP systems
There is a growing body of literature on biases in NLP systems, which has been summarized in a number of surveys (Blodgett et al., 2020; Stańczak and Augenstein, 2021; Bansal, 2022). In particular, Blodgett et al. (2020) provide a critical overview of bias in NLP systems, encompassing an analysis of 146 papers that study bias in various NLP tasks. They find that motivations for studying bias are often vague and inconsistent, lacking normative reasoning. For example, studies on embedding spaces (Bolukbasi et al., 2016; Caliskan et al., 2017; Gonen and Goldberg, 2019) illustrate biases but often do not critically engage with what constitutes bias. Similarly, works on language modeling (Lu et al., 2020; Bordia and Bowman, 2019), coreference resolution (Rudinger et al., 2018; Zhao et al., 2018), machine translation (Vanmassenhove et al., 2018; Stanovsky et al., 2019), and sentiment analysis (Kiritchenko and Mohammad, 2018) highlight bias in NLP but vary widely in their conceptualizations and methodologies. The survey recommends a deeper engagement with the relationships between language and social hierarchies and emphasizes the need for explicit statements of why certain system behaviors are harmful and to whom.
Recent advances in addressing biases in LLMs include various techniques focusing on detecting biases in the embeddings, prompt-based interactions, and contextual adaptations (Gallegos et al., 2024). However, we are not aware of any works aimed at the detection of bias in the training corpora, which is our focus.
Bias quantification approaches proposed to date have addressed three main areas: metrics for bias evaluation, datasets for bias evaluation, and methods leveraging generated text (Gallegos et al., 2024). Embedding-based metrics, such as the Sentence Encoder Association Test (May et al., 2019) and Contextualized Embedding Association Test (Guo and Caliskan, 2021), extend traditional bias metrics to contextualized embeddings. Probability-based metrics like pseudo-log-likelihood scores (Salazar et al., 2020) evaluate the likelihood of stereotypical versus anti-stereotypical sentences, while prompt-based metrics, including co-occurrence bias scores and sentiment analysis, measure bias in the generated text by comparing the distribution of terms associated with different social groups (Bordia and Bowman, 2019; Dhamala et al., 2021). These techniques provide comprehensive frameworks for assessing bias, ensuring a more nuanced and accurate understanding of how biases manifest in LLM outputs.
Biases in languages other than English
From a language perspective, most existing research on bias in NLP has focused on English. For instance, Dhamala et al. (2021) introduce the Bias in Open-Ended Language Generation Dataset (BOLD), which benchmarks social biases across five domains: profession, gender, race, religion, and political ideology, using English text generation prompts. Similarly, Webster et al. (2018) address gender bias in coreference resolution with the Gendered Ambiguous Pronouns (GAP) corpus, a gender-balanced dataset of ambiguous pronoun-name pairs.
Efforts to evaluate bias in non-English languages include the CrowS-pairs dataset, initially designed for English by Nangia et al. (2020), which provides sentence pairs to evaluate nine types of bias. Recognizing the need for multilingual approaches, Névéol et al. (2022) extended CrowS-pairs to French, creating 1,677 sentence pairs to cover stereotypes across ten types of bias, including gender and age. This extension highlighted the prevalence of biases in language models across different cultural contexts.
Other studies have explored gender bias in both English and German, revealing that methods effective for detecting bias in English may not apply as well to other languages (Bartl et al., 2020). Along those lines, a broader survey by Ramesh et al. (2023) emphasizes the need for bias evaluation and mitigation in language models beyond English, highlighting the challenges of adapting bias detection methods to multilingual contexts. This growing body of research underscores the importance of developing and refining techniques to evaluate and mitigate biases in various languages, ensuring the equitable performance of NLP systems globally.
Our contributions
While there is a body of literature on the impact of biases on various machine learning algorithms or studies related to stereotypes represented in datasets, to the best of our knowledge, there are no studies that have analyzed how to measure gender bias in texts written in gendered languages from the point of view of the prevalence of words referring to each gender. Addressing this gap is the focus of our research.
Our core contribution can be contextualized in a pyramid of methodologies to automatically detect gender bias in text corpora. Traditional word-counting methods would be located at the base of the pyramid. Unfortunately, their simplicity prevents their applicability to gendered languages like Spanish. Above this, we would find conventional NLP techniques, such as part-of-speech tagging. However, these methods might not reliably distinguish between gendered references of people versus animals or inanimate objects. At the apex of our methodological pyramid, we position our proposed LLM-based approach that not only identifies gendered words but also classifies them according to their role in conveying gender-related information within the text. Ultimately, we envision the use of these types of AI tools to automate bias evaluation with a comprehensive understanding of the underlying linguistic and social contexts.
This paper aims to shed light on the inherent challenges posed by gendered languages and propose a robust framework for detecting and measuring gender bias in Spanish text corpora. The insights derived from this study are expected to guide the development of fairer and more equitable language technologies, particularly in gendered languages.
3 Methodology
This section presents our methodology for assessing gender bias in language corpora. First, we describe a gender polarity method from the literature that has been proposed to measure gender-specific terms in English texts. We build upon this approach to propose a novel gender bias quantification method leveraging the LLMs’ natural language comprehension power to accommodate the complexities of gendered languages, using Spanish as a case study.
3.1 Gender Polarity
Most of the existing literature on assessing gender bias in language models focuses on bias quantification within the embedding space or in prompt-based interaction with an LLM. However, the scope of this paper is to measure bias in the LLM training data. The most relevant approach is presented in Dhamala et al. (2021), which introduces gender polarity as a method to quantify the presence of gender-specific language in a given text. The authors propose two metrics to evaluate gender polarity. The first one, unigram matching, involves a straightforward count of gender-specific tokens (words) from a predefined list. The tokens are defined as follows:
  • Male tokens: he, him, his, himself, man, men, he’s, boy, boys.
  • Female tokens: she, her, hers, herself, woman, women, she’s, girl, girls.
The second metric proposed in Dhamala et al. (2021) employs word embeddings to assess the proximity of words to a gendered vector space. This falls outside the scope of our work, as we focus purely on text analysis to avoid the inherent risk of amplifying biases through embeddings.
While these metrics were designed to evaluate text generation models in prompt-based interactions, specifically on the BOLD dataset (Dhamala et al., 2021), we propose extending the application of unigram matching, further referred to as the gender polarity, to quantify gender bias in larger text corpora. This approach facilitates more comprehensive assessments of gender representation in linguistic datasets.
In a given text, the number of male tokens (denoted as $G_{M}$) and the number of female tokens ($G_{F}$) are counted. The gender bias in the text can then be expressed as the ratio $G_{M}:G_{F}$. The gender polarity method was designed for English texts, where gender differentiation in language usage is mostly captured through distinct pronouns and other gender-specific tokens. Nonetheless, this approach cannot be directly applied to gendered languages, such as Spanish, by merely translating the token lists to the target language. In gendered languages, specific parts of speech, such as nouns or adjectives, have inherent gender properties that often require different forms based on the gender of the subjects or objects they describe. This leads to a more complex interaction between grammatical gender and semantic content, necessitating a tailored method capable of accounting for the intrinsic gender agreement rules in such languages.
3.2 Measuring Gender Bias in Gendered Languages
We close the gap in bias quantification for gendered languages by proposing a method that takes inspiration from the gender polarity analysis and accommodates the specific grammatical and semantic features in gendered languages. We empirically evaluate the method on Spanish as an example of a gendered language.
The following pipeline summarizes the main steps that compose the proposed framework:
1. Identification of all nouns and pronouns in a given text, which is needed to consider all potentially gendered language elements, as these are the primary carriers of gender information in Spanish.
2. Classification of each identified noun and pronoun with respect to whether it refers to a person (P) or not (N) to focus on human-related references, which reflect gender representation.
3. Determination of the grammatical gender (masculine, M; feminine, F) of each identified word, as it directly correlates to the perceived or intended gender of the person being referenced.
Note that we have consciously decided not to include adjectives, as they adapt their form to agree with the noun they modify, leading to redundancy in our analysis if those nouns are already accounted for. Furthermore, adjectives alone do not necessarily indicate the gender of the person being described unless paired directly with specific gendered nouns or pronouns, rendering the classification in step 2 impossible in many cases. Therefore, including adjectives could potentially negatively impact our method’s accuracy by introducing elements more reflective of grammatical structure rather than gender representation.
This structured approach enables a comprehensive analysis of gender representation in text, differentiating between words that refer to people and other types of words, which is crucial for gendered languages.
3.3 Gender Bias Quantification Using LLMs
Implementing the pipeline described in the previous section using classical NLP methods would typically involve a combination of tools, leveraging part-of-speech tagging for step 1 and gender classification based, for instance, on word endings and dictionary lookups for step 3. However, accurately determining whether a noun or pronoun refers to a person (step 2) would require more sophisticated semantic analysis tools.
Given these challenges, we propose to leverage LLMs for their proficiency in understanding natural language nuances and context. LLMs can seamlessly integrate all steps of the proposed pipeline. After a process of manual interactive prompt-tuning, we designed a prompt that instructs the LLM to perform noun and pronoun identification, determine if these refer to human beings, and classify their grammatical gender, all in a single query. This approach utilizes the LLM’s ability to parse and interpret complex language structures and perform multiple tasks simultaneously.
The English translation of our prompt, originally crafted in Spanish (see Appendix A), is as follows:
<EXAMPLES>
Sentence: <SENTENCE>
Identify all nouns and pronouns in the given sentence. For each of them, determine whether it refers to a person (P) or not (N), and specify its grammatical gender: masculine (M) or feminine (F). Exclude surnames. Follow the format of the provided examples without adding additional text.
We adopt a few-shot prompting strategy, showing the LLM a number of example sentences along with their analyses to prime it for the task, see Figure 1. This methodology enhances the model’s performance by providing clear examples of the desired output format and analysis procedure. Surnames are explicitly excluded from the analysis because they do not convey gender information in Spanish.
To analyze the gender representation in a given dataset, it is parsed sentence by sentence and each sentence is evaluated using the aforementioned prompt. We introduce the notation $L_{p,g}$, which denotes the number of words identified by our LLM-based approach, where $p\in\{P,N\}$ and $g\in\{M,F\}$ are binary variables that capture whether the word refers to a person ($P$) or not ($N$) and the word’s grammatical gender ($M$ for masculine, $F$ for feminine), respectively. Analogously to the gender polarity method described in Section 3.1, the gender bias in the given text is expressed by the ratio $L_{P,M}:L_{P,F}$.
Advanced state-of-the-art LLMs are used to perform this task, specifically various versions of OpenAI’s ChatGPT. These models have demonstrated significant capabilities in solving complex tasks involving nuanced language understanding and have been trained on diverse linguistic datasets. Given the intricacies of gendered languages and the complexity of the task, more advanced models from the GPT-4 family offer substantial advantages in accuracy and reliability over smaller or older model architectures.
4 Experiments
This section describes the experimental setup used to validate our methodology and to assess the extent of gender bias across various benchmark datasets. We utilize bilingual corpora to report the results using our LLM-based approach to quantify gender bias in Spanish alongside the gender polarity method for English texts.
Table 1: Performance of different LLMs on the validation dataset. The values correspond to the means and standard deviations calculated from five repetitions of the evaluation for each model.
Model Accuracy (%) Precision (%) Recall (%) F-score (%)
gpt-4-turbo-preview $85.68\pm 0.93$ $87.51\pm 0.49$ $86.58\pm 0.90$ $87.04\pm 0.61$
gpt-4o $87.57\pm 1.21$ $80.45\pm 1.35$ $89.31\pm 1.19$ $84.65\pm 1.26$
gpt-4-turbo $\mathbf{89.40\pm 0.98}$ $\mathbf{89.53\pm 0.56}$ $\mathbf{90.96\pm 0.72}$ $\mathbf{90.24\pm 0.55}$
4.1 Datasets
We evaluate the proposed method on several parallel datasets from the OPUS Machine Translation project dataset collection [^1] The website specifies that the data provided on the platform are believed to be free to redistribute. (Tiedemann, 2012) that includes aligned translations between Spanish and English. We focus on parallel datasets to facilitate the comparison of the gender polarity method for English with our LLM-based gender bias quantification method for Spanish. The datasets were chosen based on their extended use in machine translation and NLP model training in general. In particular, we evaluated our approach on the Europarl, CCAligned, Global Voices, and WMT-News datasets.
Europarl:
The Europarl dataset (Koehn, 2005) is a multilingual corpus extracted from the proceedings of the European Parliament. It is widely used for research in the NLP community, particularly for tasks such as machine translation, language modeling, and cross-lingual studies. It includes transcripts in 21 European languages, including Spanish and English. The texts cover parliamentary debates and speeches, offering a diverse range of topics and formal language styles. We use the bilingual Spanish-English dataset in version v7, covering the period from April 1996 to November 2011, which consists of 1,965,734 sentences for each language.
CCAligned:
This dataset (El-Kishky et al., 2020) is a large-scale multilingual corpus of billions of sentences derived from web-crawled data. Specifically, the corpus used in our experiments was created from 68 Common Crawl snapshots up until March 2020 (v1). It is designed to support research in tasks involving machine translation and cross-lingual studies. The bilingual Spanish-English corpus contains 15,250,346 sentences for each language.
Global Voices:
The Global Voices dataset (Nguyen and Daumé III, 2019) is a multilingual corpus collected from the Global Voices website, which features news articles and stories written by a global network of authors and translated by volunteers into multiple languages. This dataset, consisting of 359,002 sentence pairs in Spanish and English, is broadly used for research in machine translation, multilingual text analysis, and other NLP tasks. We used the version v2018q4.
WMT-News:
The WMT-News dataset is a collection of parallel corpora specifically curated for machine translation and related NLP tasks. It is associated with the annual Conference on Machine Translation (WMT, originally Workshop on Machine Translation), which organizes shared tasks to benchmark and advance the state of the art in machine translation. The version we used, v2019, contains 14,522 Spanish-English sentence pairs.
We created two representative subsets of 1,000 randomly selected sentence pairs from each of these datasets. The choice of this sample subset size was motivated by the lower bound calculated using the formula from Daniel and Cross (2018): $n=z^{2}p(1-p)/e^{2}$, where $z=2.576$ for a confidence level $\alpha=0.99$, margin of error $e=0.05$, and population proportion $p=0.5$ (worst-case scenario), yielding $n=664$. Note that this sample size is aligned with the sample size used in various NLP studies (Kreutzer et al., 2022; Ranathunga et al., 2024). Using two subsets of each dataset helps to balance achieving a certain level of generalization of the results and managing the computational costs involved.
4.2 Validation
To first validate the proposed method, the author team created a manually annotated dataset consisting of 100 sentences in Spanish extracted from the Europarl dataset. This validation dataset serves as the ground truth for assessing and comparing the performance of different LLMs and their variants, and it also allows for efficiently measurable prompt engineering.
We evaluated the following LLMs available through the OpenAI API: gpt-4-turbo-preview (gpt-4-0125-preview), gpt-4o (gpt-4o-2024-05-13), and gpt-4-turbo (gpt-4-turbo-2024-04-09). For each of them, we executed five runs of our method with the same (default) configuration. For few-shot prompting, we used five sentence-analysis pairs listed in Appendix B. The budget spent on these experiments was approximately 10 USD.
To evaluate the correctness of the LLM output, we employed a case-insensitive comparison of the identified words and the (mis)match of the two attributes ($p$ and $g$, see Section 3.3) w.r.t. the ground truth. We report the number of words that were correctly identified and correctly classified in both attributes ($n_{c}$), correctly identified but incorrectly classified in at least one attribute ($n_{i}$), missed (not identified) by the method ($n_{m}$), and extra words that do not appear in the ground truth but were returned by the method ($n_{e}$). Using these values, we define the following performance metrics:
  • Accuracy: $A=n_{c}/(n_{c}+n_{i}+n_{m})$,
  • Precision: $P=n_{c}/(n_{c}+n_{i}+n_{e})$,
  • Recall: $R=n_{c}/(n_{c}+n_{m})$,
  • F-score: $F=2PR/(P+R)$.
As reflected in Table 1, gpt-4-turbo performed the best across all metrics. Hence, we selected gpt-4-turbo as the LLM of choice for our empirical study. In addition, we tested four open-source models from the Ollama [^2] https://github.com/ollama/ollama project that can be run for inference locally: Llama 2 Uncensored (7B), Mistral (7B), Neural Chat (7B), and Llama 2 (13B). Unfortunately, all of them were found unsuitable due to their poor performance. This task appears to be too complex for these smaller models, as they did not even output the analysis in the specified format in the vast majority of cases.
4.3 Results
First, Table 2 summarizes the results of applying the gender polarity method described in Section 3.1 to the English sentences in the sample subsets of the four benchmark datasets listed in Section 4.1.
Table 2: Evaluation of two representative samples for each of the four benchmark datasets in English using the gender polarity method. The reported results follow the notation introduced in Section 3.1. The gender bias (male:female) is reported in the last column.
Dataset $G_{M}$ $G_{F}$ $G_{M}:G_{F}$
Europarl 1 32 23 1.39 : 1
Europarl 2 38 26 1.46 : 1
CCAligned 1 16 15 1.07 : 1
CCAligned 2 15 14 1.07 : 1
Global Voices 1 136 95 1.43 : 1
Global Voices 2 129 90 1.43 : 1
WMT-News 1 200 65 3.08 : 1
WMT-News 2 248 72 3.44 : 1
To evaluate the Spanish version of the datasets, we employed the gpt-4-turbo model (gpt-4-turbo-2024-04-09) with default configuration. For few-shot prompting, we used five sentence-analysis pairs listed in Appendix B. The results of the proposed method (see Section 3.3) on two representative samples of the four benchmark datasets are summarized in Table 3. The budget spent on these experiments was approximately 70 USD.
Table 3: Evaluation of two representative samples for each of the four benchmark datasets in Spanish using the proposed LLM-based method. The reported results follow the notation introduced in Section 3.3, and the asterisk denotes any value of the binary variable. The gender bias (male:female) is reported in the last column.
Dataset $L_{*,M}$ $L_{*,F}$ $L_{N,*}$ $L_{P,*}$ $L_{P,M}$ $L_{P,F}$ $L_{P,M}:L_{P,F}$
Europarl 1 3531 3131 5989 677 541 136 3.98 : 1
Europarl 2 3400 3096 5765 736 587 149 3.94 : 1
CCAligned 1 2218 1478 3388 307 246 61 4.03 : 1
CCAligned 2 2184 1510 3385 310 254 56 4.54 : 1
Global Voices 1 3205 2350 4495 1063 869 194 4.48 : 1
Global Voices 2 3237 2292 4513 1019 830 189 4.39 : 1
WMT-News 1 3576 2489 5140 929 797 132 6.04 : 1
WMT-News 2 3710 2514 5223 1001 840 161 5.22 : 1
While the male:female ratio varies significantly across datasets, all datasets are biased towards masculine references in both languages. The overall agreement of the results for the two subsets of each dataset suggests that the subsets are well-representative. The results indicate that the gender bias is substantially more pronounced in Spanish than in English, although this comparison needs to be treated with caution as the methods are different.
Interestingly, in Spanish, the lowest male:female disparity is found in the Europarl dataset, which makes intuitive sense as the European Parliament is a political institution that strives for gender equality and representation across its member states, reflecting broader societal efforts towards balanced gender representation in government and public discourse. Conversely, the largest male:female disparity is identified in the WMT-News dataset, suggesting that news reporting and journalism – especially in certain regions or topics – may still be male-dominated, reflecting broader societal trends and biases in media representation and participation.
These insights highlight the importance of using diverse data sources to train and evaluate NLP models to ensure that they do not perpetuate existing biases. The results confirm the utility of our LLM-based method in capturing and quantifying gender biases effectively.
5 Discussion
The results of our study have significant implications for the field of natural language processing, particularly in the understanding and mitigation of gender bias in gendered languages. By employing state-of-the-art LLMs to analyze Spanish corpora, our research highlights the extent to which gender bias can permeate linguistic datasets and subsequently influence machine learning models trained on these datasets. Below, we summarize the main findings of our research.
1. LLMs are a promising tool to identify gender bias in gendered corpora
There is a lack of methods to automatically detect gender disparity in text corpora in gendered languages, such as Spanish. To the best of our knowledge, we are the first to propose and empirically demonstrate the value of LLMs, especially the larger and more advanced versions like gpt-4-turbo, as a tool to effectively identify and classify gendered language within a complex linguistic framework like Spanish. This capability is crucial for developing fairer NLP applications, as it allows for a deeper understanding of how gender is represented in text and how these representations can lead to biased outcomes.
2. Gender bias pervades commonly used corpora in Spanish
We identify a significant male:female gender disparity, ranging between 4:1 (Europarl dataset) and 5–6:1 (WMT-News dataset), in commonly used benchmark datasets in Spanish. The Europarl dataset exhibits the smallest gender disparity, which is aligned with the institution’s commitment to gender equality and balanced representation. This finding suggests that legislative efforts and policies aided at promoting gender balance can significantly influence the language and content of official records. However, the number of references to males is four times the number of references to females, indicating that more efforts are needed to achieve representative gender equality. Conversely, we identify the largest gender disparity in the WMT-News dataset, which suggests that news reporting and journalism remain male-driven, reflecting broader societal trends and editorial biases in the media industry.
The prevalence of masculine w.r.t. feminine references across different datasets underscores the pervasive nature of gender bias in text corpora and the importance of identifying and mitigating such biases to create more balanced datasets, which in turn would enable more equitable models.
3. Disparities in gender representation in English vs. Spanish
The differences in gender ratio between the English and Spanish versions of our datasets further underline the linguistic and cultural variances affecting gender representation. While both languages exhibited biases, the Spanish datasets generally showed a higher degree of male dominance. This may be partly due to the grammatical structure of Spanish, which includes gendered nouns and more explicit gender marking, potentially amplifying the visibility and perpetuation of gender biases. The English datasets, while also biased, tended to have lower ratios of male:female references, possibly due to a combination of linguistic factors and different societal norms influencing the text. Interestingly, in the English corpora, the CCAligned samples exhibit the lowest gender disparity. Conversely and aligned with the findings in the Spanish case, the WMT-News samples have the largest male:female ratio. These results highlight the need for considering both linguistic and cultural contexts when analyzing and mitigating biases in multilingual and multicultural corpora.
6 Conclusion
We have presented a novel methodology for measuring gender bias in Spanish text corpora using large language models. Through experiments with four benchmark datasets, we have demonstrated the efficacy of this approach, which leverages the capabilities of LLMs to differentiate between gendered references in a nuanced manner.
Our findings reveal significant gender disparities that range between 4:1 and 6:1 in the male-to-female ratio, depending on the dataset, highlighting the need for continuous monitoring and correction of biases in linguistic data sources. The methodology developed herein not only contributes to the field of bias research in NLP but also sets a precedent for future studies aiming to understand and rectify gender bias in languages other than English, broadening the scope of bias detection and mitigation strategies to a global context. Ensuring that language technologies do not perpetuate or even amplify existing social biases is essential.
In future work, we aim to further refine our approach and explore the integration of additional linguistic and cultural nuances that may affect the representation of gender in text. Moreover, extending this framework to other gendered languages could provide a more comprehensive global perspective on gender bias in NLP. Continued advancements in model capabilities and a growing understanding of bias will undoubtedly enhance the fairness and accuracy of language-based AI systems.
Acknowledgments
The work of E.D., S.S., and N.O. on this paper was supported by a nominal grant received at the ELLIS Unit Alicante Foundation from the Regional Government of Valencia in Spain (Convenio Singular signed with Generalitat Valenciana, Conselleria de Innovación, Industria, Comercio y Turismo, Dirección General de Innovación). E.D. is also supported by Intel Corporation.
The work of Y.G. and P.M. on this paper was partially funded by the University of Alicante, the Spanish Ministry of Science and Innovation, the Generalitat Valenciana, and the European Regional Development Fund (ERDF) through the following funding: As main, VIVES: “Pla de Tecnologies de la Llengua per al valencià” project (2022/TL22/00215334) from the Projecte Estratègic per a la Recuperació i Transformació Econòmica (PERTE) and the following supportive projects: NL4DISMIS (CIPROM/2021/21) with funds granted by the Generalitat Valenciana (Conselleria d’Educacio, Investigacio, Cultura i Esport), COOLANG (PID2021-122263OB-C22), and CORTEX (PID2021-123956OB-I00) with funds granted by MCIN/AEI/10.13039/501100011033 and, as appropriate, by “ERDF A way of making Europe”, by the “European Union” or by the “European Union NextGenerationEU/PRTR”.
Limitations
While we believe that our study provides valuable insights into the measurement of gender bias in gendered languages using LLMs, we have identified several limitations.
Inclusion of epicene words: In our analysis, we included epicene words – nouns that have the same form, including grammatical gender, for male and for female referents. For example, ‘la persona’ (person) is grammatically feminine but can refer to individuals of any gender. This inherent ambiguity can lead to inaccuracies in gender classification. In the data analyzed in this work, epicene words represent 5.8 % of all identified person-referencing words. It is noteworthy that 73.7 % of epicene words were counted as feminine (see Appendix C), such that excluding the epicene words would make the gender ratios even more biased towards the masculine gender.
Contextuality of the analysis: Our methodology focuses primarily on the occurrence and classification of gendered nouns and pronouns without deeply analyzing the context in which these words appear. This approach provides a quantitative measure of gender representation but lacks the nuance of qualitative analysis. Understanding the context could reveal more about the nature of gender bias, such as whether certain roles or characteristics are predominantly associated with a specific gender. Future work should incorporate contextual analysis to provide a richer and more comprehensive understanding of gender bias in texts.
Binary gender: The current framework represents gender as a binary variable – masculine and feminine – and does not account for non-binary or other gender identities. This limitation reflects the grammatical structure of Spanish but is not aligned with the broader spectrum of gender identities present in the society. As language evolves and becomes more inclusive, future studies should consider expanding the classification system to encompass a wider range of gender identities, ensuring that NLP models are sensitive to and representative of all individuals.
Budget implications: Using commercial, state-of-the-art LLMs in research can have a non-negligible financial cost. While we selected representative subsets from widely used benchmark datasets, a more extensive analysis could potentially provide more robust and generalizable findings at a cost.
LLM dependence and model variability: Our method heavily relies on the capabilities of advanced LLMs, specifically gpt-4-turbo. The performance and accuracy of our bias detection framework are directly tied to the model’s proficiency in understanding and interpreting language. Different versions or configurations of LLMs might yield varying results, and smaller open-source models that allow for local inference currently lack the performance needed for such complex tasks. Additionally, the use of proprietary models may impact the reproducibility and transparency of our findings.
Cultural and linguistic variations: Gender representation in language can be influenced by cultural and regional variations, which our study does not explicitly address. The datasets used, while diverse, may not fully capture the nuances of different Spanish-speaking communities. Future research should consider the impact of cultural and regional differences on gender bias to develop more universally applicable methodologies.
By addressing these limitations in future work, we aim to refine our methodology, incorporate a broader range of linguistic and cultural nuances, and enhance the fairness and accuracy of gender bias detection in NLP systems.
Ethics Statement
We aim to promote fairness and inclusivity by identifying and quantifying biases in text corpora used to train LLMs. We have adhered to ethical standards and code of conduct by ensuring transparency, reproducibility, and validating our methodology against manually annotated data. The datasets used for evaluation are publicly available benchmark datasets. While our work highlights significant gender disparities, we recognize the limitations of focusing on grammar-based gender classification and the reliance on specific LLMs.
We are committed to ethical AI use, advocating for continuous improvement in bias detection and mitigation techniques. Our findings underscore the pervasive nature of gender bias in linguistic datasets and aim to inspire further research and action within the NLP community to develop more equitable language technologies. By raising awareness and providing robust methodologies, we contribute to the broader discourse on AI ethics and the pursuit of technologies that uphold human dignity and equality.
Appendix A Appendix: Prompt Formulation
Through manual interactive testing, we crafted the following prompt in Spanish, which is used in all experiments with the proposed LLM-based method reported in this paper:
<EXAMPLES>
Frase: <SENTENCE>
Instrucciones: Identifica todos los sustantivos y pronombres en la frase proporcionada. Para cada uno, determina si se refiere a un ser humano (S) o no (N), y especifica su género gramatical: masculino (M) o femenino (F). Excluye los apellidos. Sigue el formato de los ejemplos proporcionados sin añadir texto adicional.
The placeholder <EXAMPLES> is replaced with priming examples, listed in Table 5. Each of them is prepended with ‘Ejemplo #:’ (Spanish for ‘example’), where # is replaced with the example index. The placeholder <SENTENCE> is replaced with the sentence to be analyzed.
Appendix B Appendix: Few-Shot Prompting Examples
Through interactive experimenting with the LLMs, and following the common best practices, we concluded that it is beneficial to employ the few-shot prompting technique. We selected five sentences from the Europarl dataset (see Section 4.1) and provided the ground truth analysis (created manually) to prime the LLM for the bias quantification task, see Table 5.
Appendix C Appendix: Epicene Words
The proposed method counts epicene words based on their grammatical gender, although these words may refer to a person of any gender. Table 4 lists epicene words identified across all datasets analyzed in this work. Epicene words represent 5.8 % of all identified words referring to a person. The frequency analysis indicates that 258 epicene words were counted towards the feminine gender, and only 92 words were counted towards the masculine gender.
Table 4: Epicene words and their frequencies, identified across all datasets evaluated in this work using the proposed LLM-based method. Note that the word ‘miembro’ appears twice because it can be identified as feminine in specific contexts (indicated by the article ‘la’), although it generally has the masculine grammatical gender.
Word $p$ $g$ Frequency
personas $P$ $F$ 149
miembros $P$ $M$ 63
gente $P$ $F$ 54
persona $P$ $F$ 34
miembro $P$ $M$ 20
víctimas $P$ $F$ 14
individuo $P$ $M$ 7
víctima $P$ $F$ 5
miembro $P$ $F$ 2
individuos $P$ $M$ 2
Table 5: Few-shot prompting examples used in the experiments. Note that although we indicate person-referencing words with ‘P’ in the paper, the list of words employed in the prompt in Spanish uses ‘S’ instead.
Sentence Analysis
El señor Presidente viajó a Tokio para reunirse con el secretario de estado y a la mañana siguiente tuvo que volar a Madrid por temas personales.
señor – S, M
Presidente – S, M
Tokio – N, M
secretario – S, M
estado – N, M
mañana – N, F
Madrid – N, M
temas – N, M
Mi colega Sr. Allan Hofmann se dirigió a los ciudadanos de Madrid, recordándoles que son personas con derechos y responsabilidades.
colega – S, M
Sr. – S, M
Allan – S, M
ciudadanos – S, M
Madrid – N, M
personas – S, F
derechos – N, M
responsabilidades – N, F
El señor Presidente de la comisión de educación se reunió con los estudiantes en Tokio, donde el distinguido Sir Ben Smith compartió su visión sobre el futuro de la enseñanza.
señor – S, M
Presidente – S, M
comisión – N, F
educación – N, F
estudiantes – S, M
Tokio – N, M
Sir – S, M
Ben – S, M
visión – N, F
futuro – N, M
enseñanza – N, F
El Sr. Johnson, un respetado colega de la ciudadanía británica, ha vivido en Londres durante más de dos décadas, donde trabaja incansablemente para mejorar la comunidad local.
Sr. – S, M
colega – S, M
ciudadanía – N, F
Londres – N, M
décadas – N, F
comunidad – N, F
Encontré en Europa no solo destinos turísticos, sino un hogar temporal donde me sentí ciudadana del mundo, abrazando la diversidad y la riqueza cultural que esta tierra ofrece.
Europa – N, F
destinos – N, M
hogar – N, M
ciudadana – S, F
mundo – N, M
diversidad – N, F
riqueza – N, F
tierra – N, F

### v2

@@
-Leveraging Large Language Models to Measure Gender Bias in Gendered Languages
+Leveraging Large Language Models to Measure Gender Representation Bias in Gendered Language Corpora
 Abstract
-Gender bias in text corpora used in various natural language processing (NLP) contexts, such as for training large language models (LLMs), can lead to the perpetuation and amplification of societal inequalities.
-This is particularly pronounced in gendered languages like Spanish or French, where grammatical structures inherently encode gender, making the bias analysis more challenging.
-Existing methods designed for English are inadequate for this task due to the intrinsic linguistic differences between English and gendered languages.
-This paper introduces a novel methodology that leverages the contextual understanding capabilities of LLMs to quantitatively analyze gender representation in Spanish corpora.
-By utilizing LLMs to identify and classify gendered nouns and pronouns in relation to their reference to human entities, our approach provides a nuanced analysis of gender biases.
-We empirically validate our method on four widely-used benchmark datasets, uncovering significant gender disparities with a male-to-female ratio ranging from 4:1 to 6:1.
-These findings demonstrate the value of our methodology for bias quantification in gendered languages and suggest its application in NLP, contributing to the development of more equitable language technologies.
-Leveraging Large Language Models to Measure
-Gender Bias in Gendered Languages
+Gender bias in text corpora that are used for a variety of natural language processing (NLP) tasks, such as for training large language models (LLMs), can lead to the perpetuation and amplification of societal inequalities.
+This phenomenon is particularly pronounced in gendered languages like Spanish or French, where grammatical structures inherently encode gender, making the bias analysis more challenging.
+A first step in quantifying gender bias in text entails computing biases in gender representation, i.e., differences in the prevalence of words referring to males vs.
+females.
+Existing methods to measure gender representation bias in text corpora have mainly been proposed for English and do not generalize to gendered languages due to the intrinsic linguistic differences between English and gendered languages.
+This paper introduces a novel methodology that leverages the contextual understanding capabilities of LLMs to quantitatively measure gender representation bias in Spanish corpora.
+By utilizing LLMs to identify and classify gendered nouns and pronouns in relation to their reference to human entities, our approach provides a robust analysis of gender representation bias in gendered languages.
+We empirically validate our method on four widely-used benchmark datasets, uncovering significant gender prevalence disparities with a male-to-female ratio ranging from 4:1 to 6:1.
+These findings demonstrate the value of our methodology for bias quantification in gendered language corpora and suggest its application in NLP, contributing to the development of more equitable language technologies.
+Leveraging Large Language Models to Measure Gender Representation Bias in Gendered Language Corpora
 Erik Derner1, Sara Sansalvador de la Fuente1, Yoan Gutiérrez2, Paloma Moreda2, Nuria Oliver1 1 ELLIS Alicante, Alicante, Spain  2 University of Alicante, Alicante, Spain Correspondence: erik@ellisalicante.org
 1 Introduction
-Figure 1: Overview of the proposed method for gender bias detection and measurement in gendered languages using LLMs.
+Figure 1: Overview of the proposed method for the detection and measurement of biases in the prevalence of genders in gendered language corpora using LLMs.
 In recent years, the presence of social biases in machine learning models (Barocas et al., 2019) has gained significant attention due to their potential to perpetuate and amplify existing inequalities, impacting areas of great consequence in people’s lives, such as hiring practices (Raghavan et al., 2020), law enforcement (Babuta and Oswald, 2019), healthcare (Panch et al., 2019), and everyday digital interactions.
 Among various forms of bias, gender bias is particularly concerning because it affects roughly half of the global population and has pervasive effects across different sectors of society.
-This concern is amplified in the area of natural language processing (NLP), where biases present in training corpora can propagate through machine learning models and impact a broad range of applications.
-In this work, we consider bias in text corpora used to train large language models (LLMs).
-The main motivation for studying bias in training data of LLMs stems from the lifecycle of these models.
+This concern is amplified in the area of natural language processing (NLP), where biases present in training corpora can propagate through machine learning models and impact a broad range of applications (Kiritchenko and Mohammad, 2018; Costa-jussà and de Jorge, 2020).
+This paper focuses on representation bias – a type of bias where certain genders are underrepresented, stereotyped or portrayed in a limited manner – in the language corpora that are used to train large language models (LLMs).
+A core motivation for our study stems from the lifecycle of LLMs.
 From training to deployment, biases in the training data can not only be perpetuated but even amplified by the models (Kotek et al., 2023).
-This data often comes from diverse sources, including books, websites, social media, and other digital texts.
+The training corpora typically is obtained from diverse sources, including books, websites, social media, and other digital texts.
 If these sources contain biases – which they frequently do, reflecting societal prejudices and stereotypes – the models learn these biases as part of their training process.
-Hence, the first stage – data curation – becomes of paramount importance.
+Hence, the first stage of data curation is of paramount importance.
 It involves selecting, cleaning, and annotating the training data to minimize biases as much as possible.
-Effective data curation requires a nuanced understanding of the types of biases that can exist and the methods to mitigate them, including removing or balancing biased content, ensuring diverse and representative data samples, and incorporating fairness and ethical guidelines in the data selection process (Gallegos et al., 2024).
-However, existing methods for evaluating bias in text, such as gender polarity (Dhamala et al., 2021) and co-occurrences (Bordia and Bowman, 2019), have been primarily developed for English.
-These approaches often rely on the neutrality of the English language in terms of grammatical gender, allowing for relatively straightforward identification and categorization of gendered expressions.
-Nonetheless, such an approach is not directly applicable to gendered languages like Spanish, where nouns, articles, adjectives and other word forms reflect gender, making the detection and analysis of gender bias a much harder problem to solve.
-Given the limitations of existing methods and the prevalence of gendered languages in the world, the focus of this paper is to develop a novel and robust method to detect gender bias in corpora used to train LLMs in Spanish, a representative gendered language spoken by over 500 million people across diverse cultures and regions in the world.
-In Spanish, every noun carries a gender marker but only the words that refer to human beings are reflective of a potential gender bias.
-We propose a novel methodology that leverages state-of-the-art LLMs to detect and analyze gender bias in Spanish text corpora.
+Effective data curation requires an understanding of the types of biases that can exist and the methods to mitigate them, including removing or balancing biased content, ensuring diverse and representative data samples, and incorporating fairness and ethical guidelines in the data selection process (Gallegos et al., 2024).
+In languages with grammatical gender (gendered languages), such as Spanish or French, this issue becomes even more complex.
+The grammatical structure may invisibly amplify certain gendered associations, impacting how often and in what context different genders are represented.
+In fact, research has shown that languages with grammatical gender exhibit more explicit gender biases due to frequent gender marking, which differs substantially from English, where gendered language use is mostly limited to pronouns and a few gender-specific terms (Kiritchenko and Mohammad, 2018; Costa-jussà and de Jorge, 2020).
+For example, masculine terms are often considered the ‘default’ in many gendered languages, which can implicitly prioritize male entities or perspectives.
+Studies indicate that grammatical gender affects model behavior in downstream tasks, including machine translation and sentiment analysis, leading to skewed model predictions that can disadvantage one gender over another (Gonen et al., 2019; Omrani Sabbaghi and Caliskan, 2022).
+For instance, Costa-jussà and de Jorge (Costa-jussà and de Jorge, 2020) demonstrate that machine translation systems tend to propagate gender biases more overtly in gendered languages.
+Consequently, adapting bias measurement methods to account for language-specific characteristics is essential for accurate assessments, as traditional methods used for English may not fully capture the nuances of gender bias in gendered languages like Spanish (Hellinger and Bußmann, 2001; Cho et al., 2021).
+Given the limitations of existing methods and the prevalence of gendered languages in the world, the focus of this paper is to develop a novel and robust method to detect gender representation bias in corpora used to train LLMs in Spanish, a representative gendered language spoken by over 500 million people across diverse cultures and regions in the world.
+In Spanish, every noun carries a gender marker but only the words that refer to human beings are reflective of a potential gender representation bias.
+To this end, we propose a novel methodology that leverages state-of-the-art LLMs to detect and analyze gender representation bias in Spanish text corpora.
 Unlike traditional methods that rely on part-of-speech tagging or specific machine learning models to detect person references, our approach benefits from the contextual understanding capabilities of LLMs.
 By prompting LLMs to identify and classify nouns and pronouns by their reference to persons and their grammatical gender, we introduce a method that respects the linguistic structure of Spanish while providing the necessary depth of analysis.
+Our methodology offers a proactive approach, enabling the identification of gender biases at the data level rather than addressing them post-training.
+By focusing on the unique grammatical structures of Spanish, we aim to provide a tool to help detect and mitigate biases early, contributing to the development of more equitable language models.
 We first provide an overview of the most relevant literature on biases in NLP systems in English and non-English languages, followed by a detailed description of the proposed methodology.
@@
 Interestingly, the English translations of the same corpora exhibit a male:female gender ratio of 1:1 to 3.5:1.
-Our results illustrate the importance of developing novel methods to detect gender bias in text corpora in gendered languages, such as Spanish.
+Our results illustrate the importance of developing novel methods to detect gender representation bias in text corpora in gendered languages, such as Spanish.
 We conclude with a discussion, limitations, ethical implications, and lines of future work.
@@
 They find that motivations for studying bias are often vague and inconsistent, lacking normative reasoning.
-For example, studies on embedding spaces (Bolukbasi et al., 2016; Caliskan et al., 2017; Gonen and Goldberg, 2019) illustrate biases but often do not critically engage with what constitutes bias.
-Similarly, works on language modeling (Lu et al., 2020; Bordia and Bowman, 2019), coreference resolution (Rudinger et al., 2018; Zhao et al., 2018), machine translation (Vanmassenhove et al., 2018; Stanovsky et al., 2019), and sentiment analysis (Kiritchenko and Mohammad, 2018) highlight bias in NLP but vary widely in their conceptualizations and methodologies.
+Works on language modeling (Lu et al., 2020; Bordia and Bowman, 2019), coreference resolution (Rudinger et al., 2018; Zhao et al., 2018), machine translation (Vanmassenhove et al., 2018; Stanovsky et al., 2019), and sentiment analysis (Kiritchenko and Mohammad, 2018) highlight bias in NLP but vary widely in their conceptualizations and methodologies.
 The survey recommends a deeper engagement with the relationships between language and social hierarchies and emphasizes the need for explicit statements of why certain system behaviors are harmful and to whom.
+Gender imbalances in the training corpora are a significant and well-documented source of bias in machine learning.
+Numerous studies have shown that these imbalances can significantly influence model behavior.
+For example, Bolukbasi et al.
+(2016) found that word embeddings trained on biased data can amplify gender stereotypes in tasks like word similarity and analogical reasoning.
+Similarly, studies by Caliskan et al.
+(2017) and Brunet et al.
+(2019) demonstrate that biases present in training corpora can directly influence model outputs, perpetuating gendered stereotypes and imbalances in downstream tasks.
+Therefore, detecting and addressing gender imbalances in corpora is crucial to mitigating bias effects.
 Recent advances in addressing biases in LLMs include various techniques focusing on detecting biases in the embeddings, prompt-based interactions, and contextual adaptations (Gallegos et al., 2024).
-However, we are not aware of any works aimed at the detection of bias in the training corpora, which is our focus.
+However, we are not aware of any works aimed at the detection of gender representation bias in the training corpora, which is our focus.
 Bias quantification approaches proposed to date have addressed three main areas: metrics for bias evaluation, datasets for bias evaluation, and methods leveraging generated text (Gallegos et al., 2024).
@@
 Probability-based metrics like pseudo-log-likelihood scores (Salazar et al., 2020) evaluate the likelihood of stereotypical versus anti-stereotypical sentences, while prompt-based metrics, including co-occurrence bias scores and sentiment analysis, measure bias in the generated text by comparing the distribution of terms associated with different social groups (Bordia and Bowman, 2019; Dhamala et al., 2021).
-These techniques provide comprehensive frameworks for assessing bias, ensuring a more nuanced and accurate understanding of how biases manifest in LLM outputs.
+These techniques provide comprehensive frameworks for assessing bias, enabling an accurate understanding of how biases manifest in LLM outputs.
 Biases in languages other than English
-From a language perspective, most existing research on bias in NLP has focused on English.
+From a language perspective, most existing research about biases in NLP has focused on English.
 For instance, Dhamala et al.
@@
 Similarly, Webster et al.
-(2018) address gender bias in coreference resolution with the Gendered Ambiguous Pronouns (GAP) corpus, a gender-balanced dataset of ambiguous pronoun-name pairs.
+(2018) address gender bias in co-reference resolution with the Gendered Ambiguous Pronouns (GAP) corpus, a gender-balanced dataset of ambiguous pronoun-name pairs.
 Efforts to evaluate bias in non-English languages include the CrowS-pairs dataset, initially designed for English by Nangia et al.
@@
 Our contributions
-While there is a body of literature on the impact of biases on various machine learning algorithms or studies related to stereotypes represented in datasets, to the best of our knowledge, there are no studies that have analyzed how to measure gender bias in texts written in gendered languages from the point of view of the prevalence of words referring to each gender.
+While there is a body of literature on the impact of biases on various machine learning algorithms or studies related to stereotypes represented in datasets, to the best of our knowledge, there are no studies that have analyzed how to measure gender representation bias in texts written in gendered languages from the point of view of the prevalence of words referring to each gender.
 Addressing this gap is the focus of our research.
-Our core contribution can be contextualized in a pyramid of methodologies to automatically detect gender bias in text corpora.
-Traditional word-counting methods would be located at the base of the pyramid.
+Our core contribution can be contextualized in a pyramid of methodologies to automatically detect gender representation bias in text corpora.
+Traditional male-female word-counting methods would be located at the base of the pyramid.
 Unfortunately, their simplicity prevents their applicability to gendered languages like Spanish.
@@
 The insights derived from this study are expected to guide the development of fairer and more equitable language technologies, particularly in gendered languages.
+While our primary focus is on detecting and quantifying gender representation bias, these capabilities represent a crucial first step in the debiasing process.
+By providing a reliable means to identify biases, our method enables the development of targeted interventions and ultimately supports creating more inclusive NLP applications.
 3 Methodology
-This section presents our methodology for assessing gender bias in language corpora.
+This section presents our methodology for assessing gender representation bias in language corpora.
 First, we describe a gender polarity method from the literature that has been proposed to measure gender-specific terms in English texts.
-We build upon this approach to propose a novel gender bias quantification method leveraging the LLMs’ natural language comprehension power to accommodate the complexities of gendered languages, using Spanish as a case study.
+We build upon this approach to propose a novel gender representation bias quantification method leveraging the LLMs’ natural language comprehension power to accommodate the complexities of gendered languages, using Spanish as a case study.
 3.1 Gender Polarity
@@
 In a given text, the number of male tokens (denoted as $G_{M}$) and the number of female tokens ($G_{F}$) are counted.
-The gender bias in the text can then be expressed as the ratio $G_{M}:G_{F}$.
+The gender representation bias in the text can then be expressed as the ratio $G_{M}:G_{F}$.
 The gender polarity method was designed for English texts, where gender differentiation in language usage is mostly captured through distinct pronouns and other gender-specific tokens.
@@
 This leads to a more complex interaction between grammatical gender and semantic content, necessitating a tailored method capable of accounting for the intrinsic gender agreement rules in such languages.
-3.2 Measuring Gender Bias in Gendered Languages
-We close the gap in bias quantification for gendered languages by proposing a method that takes inspiration from the gender polarity analysis and accommodates the specific grammatical and semantic features in gendered languages.
+3.2 Measuring Gender Representation Bias in Gendered Languages
+We close the gap in the quantification of gender representation bias in gendered languages by proposing a method that takes inspiration from the gender polarity analysis and accommodates the specific grammatical and semantic features in gendered languages.
 We empirically evaluate the method on Spanish as an example of a gendered language.
@@
 3. Determination of the grammatical gender (masculine, M; feminine, F) of each identified word, as it directly correlates to the perceived or intended gender of the person being referenced.
-Note that we have consciously decided not to include adjectives, as they adapt their form to agree with the noun they modify, leading to redundancy in our analysis if those nouns are already accounted for.
-Furthermore, adjectives alone do not necessarily indicate the gender of the person being described unless paired directly with specific gendered nouns or pronouns, rendering the classification in step 2 impossible in many cases.
-Therefore, including adjectives could potentially negatively impact our method’s accuracy by introducing elements more reflective of grammatical structure rather than gender representation.
+Note that we have consciously decided not to include articles, determinants and adjectives, as they adapt their form to agree with the noun they modify, leading to redundancy in our analysis if those nouns are already accounted for.
 This structured approach enables a comprehensive analysis of gender representation in text, differentiating between words that refer to people and other types of words, which is crucial for gendered languages.
-3.3 Gender Bias Quantification Using LLMs
-Implementing the pipeline described in the previous section using classical NLP methods would typically involve a combination of tools, leveraging part-of-speech tagging for step 1 and gender classification based, for instance, on word endings and dictionary lookups for step 3.
+An important consideration in analyzing Spanish is the traditional convention of using the male plural form to refer to groups that may include both men and women (e.g., doctores for doctors, including both male and female doctors).
+This linguistic norm inherently assigns the male grammatical gender to such mixed-gender groups, leading our method to classify these terms as male.
+This convention, although prevalent in many gendered languages, can contribute to the under-representation of females.
+Currently, in Spanish as in other gendered languages, listing explicitly both genders is the preferred form and has become the new standard [^1] https://www.unwomen.org/sites/default/files/Headquarters/Attachments/Sections/Library/Gender-inclusive%20language/Guidelines-on-gender-inclusive-language-es.pdf (e.g., doctoras y doctores collectively referring to female and male doctors).
+Therefore, considering the generic male plural as a form of gender representation bias is justified.
+3.3 Gender Representation Bias Quantification Using LLMs
+Implementing the pipeline described in the previous section by means of classical NLP methods would typically involve a combination of tools, leveraging part-of-speech tagging for step 1 and gender classification based, for instance, on word endings and dictionary lookups for step 3.
 However, accurately determining whether a noun or pronoun refers to a person (step 2) would require more sophisticated semantic analysis tools.
-Given these challenges, we propose to leverage LLMs for their proficiency in understanding natural language nuances and context.
+Given these challenges, we propose to leverage state-of-the-art LLMs for their proficiency in understanding natural language nuances and context.
 LLMs can seamlessly integrate all steps of the proposed pipeline.
@@
 This approach utilizes the LLM’s ability to parse and interpret complex language structures and perform multiple tasks simultaneously.
+An important advantage of our method is its scalability to other gendered languages beyond Spanish.
+The use of state-of-the-art multilingual large language models enables our approach to handle a wide range of languages that incorporate grammatical gender.
+These models have been trained on diverse linguistic data, allowing them to recognize and analyze gendered terms in languages with varying structures and grammatical rules.
 The English translation of our prompt, originally crafted in Spanish (see Appendix A), is as follows:
@@
 We introduce the notation $L_{p,g}$, which denotes the number of words identified by our LLM-based approach, where $p\in\{P,N\}$ and $g\in\{M,F\}$ are binary variables that capture whether the word refers to a person ($P$) or not ($N$) and the word’s grammatical gender ($M$ for masculine, $F$ for feminine), respectively.
-Analogously to the gender polarity method described in Section 3.1, the gender bias in the given text is expressed by the ratio $L_{P,M}:L_{P,F}$.
+Analogously to the gender polarity method described in Section 3.1, the gender representation bias in the given text is expressed by the ratio $L_{P,M}:L_{P,F}$.
 Advanced state-of-the-art LLMs are used to perform this task, specifically various versions of OpenAI’s ChatGPT.
@@
 (Tiedemann, 2012) that includes aligned translations between Spanish and English.
-We focus on parallel datasets to facilitate the comparison of the gender polarity method for English with our LLM-based gender bias quantification method for Spanish.
+We focus on parallel datasets to facilitate the comparison of the gender polarity method for English with our LLM-based gender representation bias quantification method for Spanish.
 The datasets were chosen based on their extended use in machine translation and NLP model training in general.
@@
 The budget spent on these experiments was approximately 10 USD.
+Note that the objective of our work is not to compare the performance of different LLMs in detecting gender imbalances but rather to show that, with careful prompt engineering, state-of-the-art LLMs can accurately identify gender representation imbalances in corpora of gendered languages.
 To evaluate the correctness of the LLM output, we employed a case-insensitive comparison of the identified words and the (mis)match of the two attributes ($p$ and $g$, see Section 3.3) w.r.t.
 the ground truth.
-We report the number of words that were correctly identified and correctly classified in both attributes ($n_{c}$), correctly identified but incorrectly classified in at least one attribute ($n_{i}$), missed (not identified) by the method ($n_{m}$), and extra words that do not appear in the ground truth but were returned by the method ($n_{e}$).
+We compute the number of words that were correctly identified and correctly classified in both attributes ($n_{c}$), correctly identified but incorrectly classified in at least one attribute ($n_{i}$), missed (not identified) by the method ($n_{m}$), and extra words that do not appear in the ground truth but were returned by the method ($n_{e}$).
 Using these values, we define the following performance metrics:
@@
 In addition, we tested four open-source models from the Ollama [^3] https://github.com/ollama/ollama project that can be run for inference locally: Llama 2 Uncensored (7B), Mistral (7B), Neural Chat (7B), and Llama 2 (13B).
-Unfortunately, all of them were found unsuitable due to their poor performance.
-This task appears to be too complex for these smaller models, as they did not even output the analysis in the specified format in the vast majority of cases.
+Unfortunately, all of them were found unsuitable for this task due to their poor performance.
+This task appears to be too complex for these smaller models, as they did not even produce a meaningful output in the vast majority of cases.
 4.3 Results
 First, Table 2 summarizes the results of applying the gender polarity method described in Section 3.1 to the English sentences in the sample subsets of the four benchmark datasets listed in Section 4.1.
-Table 2: Evaluation of two representative samples for each of the four benchmark datasets in English using the gender polarity method.
+Table 2: Gender representation results on two representative samples for each of the four benchmark datasets in English using the gender polarity method.
 The reported results follow the notation introduced in Section 3.1.
-The gender bias (male:female) is reported in the last column.
+The gender representation bias (male:female) is reported in the last column.
 Dataset $G_{M}$ $G_{F}$ $G_{M}:G_{F}$
@@
 The budget spent on these experiments was approximately 70 USD.
-Table 3: Evaluation of two representative samples for each of the four benchmark datasets in Spanish using the proposed LLM-based method.
+Table 3: Gender representation results on two representative samples for each of the four benchmark datasets in Spanish using the proposed LLM-based method.
 The reported results follow the notation introduced in Section 3.3, and the asterisk denotes any value of the binary variable.
-The gender bias (male:female) is reported in the last column.
+The gender representation bias (male:female) is reported in the last column.
 Dataset $L_{*,M}$ $L_{*,F}$ $L_{N,*}$ $L_{P,*}$ $L_{P,M}$ $L_{P,F}$ $L_{P,M}:L_{P,F}$
@@
 The overall agreement of the results for the two subsets of each dataset suggests that the subsets are well-representative.
-The results indicate that the gender bias is substantially more pronounced in Spanish than in English, although this comparison needs to be treated with caution as the methods are different.
-Interestingly, in Spanish, the lowest male:female disparity is found in the Europarl dataset, which makes intuitive sense as the European Parliament is a political institution that strives for gender equality and representation across its member states, reflecting broader societal efforts towards balanced gender representation in government and public discourse.
-Conversely, the largest male:female disparity is identified in the WMT-News dataset, suggesting that news reporting and journalism – especially in certain regions or topics – may still be male-dominated, reflecting broader societal trends and biases in media representation and participation.
+The results indicate that the gender representation bias is substantially more pronounced in Spanish than in English, although this comparison needs to be treated with caution as the methods are different.
 These insights highlight the importance of using diverse data sources to train and evaluate NLP models to ensure that they do not perpetuate existing biases.
-The results confirm the utility of our LLM-based method in capturing and quantifying gender biases effectively.
+The results confirm the utility of our LLM-based method in capturing and quantifying gender representation bias effectively.
 5 Discussion
 The results of our study have significant implications for the field of natural language processing, particularly in the understanding and mitigation of gender bias in gendered languages.
-By employing state-of-the-art LLMs to analyze Spanish corpora, our research highlights the extent to which gender bias can permeate linguistic datasets and subsequently influence machine learning models trained on these datasets.
-Below, we summarize the main findings of our research.
-1. LLMs are a promising tool to identify gender bias in gendered corpora
-There is a lack of methods to automatically detect gender disparity in text corpora in gendered languages, such as Spanish.
+By employing state-of-the-art LLMs to analyze Spanish corpora, our research highlights the extent to which gender representation bias can permeate linguistic datasets and subsequently influence machine learning models trained on these datasets.
+Below, we provide a discussion of the main findings of our research.
+1. LLMs are a promising tool to identify gender representation bias in gendered corpora
+There is a lack of methods to automatically detect gender representation bias in text corpora in gendered languages, such as Spanish.
 To the best of our knowledge, we are the first to propose and empirically demonstrate the value of LLMs, especially the larger and more advanced versions like gpt-4-turbo, as a tool to effectively identify and classify gendered language within a complex linguistic framework like Spanish.
 This capability is crucial for developing fairer NLP applications, as it allows for a deeper understanding of how gender is represented in text and how these representations can lead to biased outcomes.
-2. Gender bias pervades commonly used corpora in Spanish
-We identify a significant male:female gender disparity, ranging between 4:1 (Europarl dataset) and 5–6:1 (WMT-News dataset), in commonly used benchmark datasets in Spanish.
-The Europarl dataset exhibits the smallest gender disparity, which is aligned with the institution’s commitment to gender equality and balanced representation.
+2. Gender representation bias pervades commonly used corpora in Spanish
+We identify a significant male:female gender representation disparity, ranging between 4:1 (Europarl dataset) and 5–6:1 (WMT-News dataset), in commonly used benchmark datasets in Spanish.
+The Europarl dataset exhibits the smallest gender representation disparity, which is aligned with the institution’s commitment to gender equality and balanced representation.
 This finding suggests that legislative efforts and policies aided at promoting gender balance can significantly influence the language and content of official records.
 However, the number of references to males is four times the number of references to females, indicating that more efforts are needed to achieve representative gender equality.
-Conversely, we identify the largest gender disparity in the WMT-News dataset, which suggests that news reporting and journalism remain male-driven, reflecting broader societal trends and editorial biases in the media industry.
+Conversely, we identify the largest gender representation disparity in the WMT-News dataset, which suggests that news reporting and journalism remain male-driven, reflecting broader societal trends and editorial biases in the media industry.
 The prevalence of masculine w.r.t.
-feminine references across different datasets underscores the pervasive nature of gender bias in text corpora and the importance of identifying and mitigating such biases to create more balanced datasets, which in turn would enable more equitable models.
+feminine references across different datasets underscores the pervasive nature of gender representation bias in text corpora and the importance of identifying and mitigating such biases to create more balanced datasets, which in turn would enable more equitable models.
 3. Disparities in gender representation in English vs. Spanish
@@
 This may be partly due to the grammatical structure of Spanish, which includes gendered nouns and more explicit gender marking, potentially amplifying the visibility and perpetuation of gender biases.
-The English datasets, while also biased, tended to have lower ratios of male:female references, possibly due to a combination of linguistic factors and different societal norms influencing the text.
-Interestingly, in the English corpora, the CCAligned samples exhibit the lowest gender disparity.
-Conversely and aligned with the findings in the Spanish case, the WMT-News samples have the largest male:female ratio.
+The English datasets, while also biased, tend to have lower ratios of male:female references, possibly due to a combination of linguistic factors and different societal norms influencing the text.
+Interestingly, in the English corpora, the CCAligned samples exhibit the lowest gender representation disparity.
+Conversely and aligned with the findings for Spanish, the WMT-News samples have the largest male:female ratio.
 These results highlight the need for considering both linguistic and cultural contexts when analyzing and mitigating biases in multilingual and multicultural corpora.
+4. Representation bias and its impact
+While considering the imbalance of gendered words in training corpora as a measure of bias can be viewed as simplifying the problem, it represents an essential first step, particularly in gendered languages, as it provides a quantitative assessment of representation bias.
+This type of bias is one of the three identified forms of gender bias in language, which include: ‘language that ignores women; language that defines women narrowly; and language that depreciates women’ (Martin, 2020; Weatherall et al., 2002).
+Our focus is on the first form – language that ignores women – by measuring the differences in the prevalence of words that refer to male versus female human entities.
+Discourse analysis has demonstrated that imbalanced portrayals or frequency of gender mentions can lead to representational biases, as language often reflects societal norms and values.
+An overrepresentation of male-oriented language implies a bias toward male perspectives, effectively underrepresenting women and making female perspectives less visible (Pauwels, 2003).
+By quantifying these imbalances, we can lay the foundation for more comprehensive bias-mitigation strategies, which can later be applied to address subtler forms of bias in model outputs and behavior (Zhao et al., 2017; Babaeianjelodar et al., 2020).
+Importantly, the main contribution of this paper is to provide a method to measure gender representation imbalances in gendered languages, filling a gap in existing research where such a tool has yet to be developed.
+Our work is framed within a larger initiative aimed at training large language models from scratch in non-English languages, where it is crucial to assess and mitigate gender biases within the training corpora (Wolfe and Caliskan, 2021).
+Biases originating in training data, particularly an imbalance in male/female words, can lead to biased models that replicate and potentially amplify these biases.
+When training data is predominantly male-oriented, models are more likely to generate outputs that favor male markers and perspectives.
+As such, addressing these biases early in the model development process, as we propose, is critical to creating balanced and equitable language models for non-English languages.
+5. Implications for debiasing strategies
+While the primary contribution of our paper is a novel method for quantifying gender representation bias in corpora, it is important to recognize that accurate detection and quantification of bias is a foundational first step in the debiasing process.
+Without reliable methods to identify and quantify bias, subsequent debiasing efforts may lack precision and effectiveness.
+Our method fills a critical gap by offering a systematic and robust approach to detect biases that are often overlooked, particularly in gendered languages.
+The insights gained from our methodology can directly inform and enhance debiasing strategies by identifying specific instances and patterns of bias in training corpora.
+For example, if our tool detects that a dataset consistently uses masculine references, such as ‘el doctor’ for ‘the doctor’, this insight enables the development of targeted interventions to mitigate the bias.
+Potential debiasing strategies that our method could support include:
+  • Gender-Neutral Rewriting: Once gender-marked words and phrases are identified, they can be rewritten in a gender-neutral manner where contextually appropriate.
+For instance, ‘el doctor’ could be replaced with ‘el personal médico’ (the medical personnel), which avoids gender marking.
+  • Balanced Representation: To balance gender representation, sentences with the opposite gender marker can be added.
+For example, if ‘el doctor’ predominates, additional sentences featuring the female version of the word, ‘la doctora’, could be incorporated to balance the dataset.
+  • Fair Data Sampling: During training, the sampling probability of gendered terms can be dynamically adjusted to maintain gender balance.
+For instance, a lower sampling rate might be applied to more frequently occurring masculine terms, which helps prevent their overrepresentation in the trained model.
 6 Conclusion
-We have presented a novel methodology for measuring gender bias in Spanish text corpora using large language models.
+We have presented a novel methodology for measuring gender representation bias in Spanish text corpora using large language models.
 Through experiments with four benchmark datasets, we have demonstrated the efficacy of this approach, which leverages the capabilities of LLMs to differentiate between gendered references in a nuanced manner.
-Our findings reveal significant gender disparities that range between 4:1 and 6:1 in the male-to-female ratio, depending on the dataset, highlighting the need for continuous monitoring and correction of biases in linguistic data sources.
+Our findings reveal significant gender representation disparities that range between 4:1 and 6:1 in the male-to-female ratio, depending on the dataset, highlighting the need for continuous monitoring and correction of biases in linguistic data sources.
 The methodology developed herein not only contributes to the field of bias research in NLP but also sets a precedent for future studies aiming to understand and rectify gender bias in languages other than English, broadening the scope of bias detection and mitigation strategies to a global context.
 Ensuring that language technologies do not perpetuate or even amplify existing social biases is essential.
-In future work, we aim to further refine our approach and explore the integration of additional linguistic and cultural nuances that may affect the representation of gender in text.
-Moreover, extending this framework to other gendered languages could provide a more comprehensive global perspective on gender bias in NLP.
-Continued advancements in model capabilities and a growing understanding of bias will undoubtedly enhance the fairness and accuracy of language-based AI systems.
+While our current method focuses specifically on detecting structural biases, it represents a foundational step in a larger framework aimed at addressing various contextual bias types.
+This focus aligns with the ‘methodological pyramid’ described at the end of Section 2, where gendered markers, though not the only type of bias, are a well-documented and impactful source of gender bias in language models.
+In future work, we aim to expand upon this foundation to incorporate additional forms of contextual bias.
+Leveraging the semantic capabilities of LLMs, we plan to refine our approach to detect more subtle and complex gender bias patterns embedded in language.
+The insights gained from our focus on structural bias can guide broader strategies for bias quantification and mitigation, supporting the development of models that capture linguistic and social contexts with greater nuance.
+This progression will ultimately enable a more comprehensive approach to fair and ethical language model development.
 Acknowledgments
@@
 E.D.
-is also supported by Intel Corporation.
+and N.O.
+are also supported by Intel Corporation.
+E.D.
+is also supported by the VIVES: ‘Pla de Tecnologies de la Llengua per al valencià’ project (2022/TL22/00215334) from the Projecte Estratègic per a la Recuperació i Transformació Econòmica (PERTE).
 The work of Y.G.
 and P.M.
-on this paper was partially funded by the University of Alicante, the Spanish Ministry of Science and Innovation, the Generalitat Valenciana, and the European Regional Development Fund (ERDF) through the following funding: As main, VIVES: “Pla de Tecnologies de la Llengua per al valencià” project (2022/TL22/00215334) from the Projecte Estratègic per a la Recuperació i Transformació Econòmica (PERTE) and the following supportive projects: NL4DISMIS (CIPROM/2021/21) with funds granted by the Generalitat Valenciana (Conselleria d’Educacio, Investigacio, Cultura i Esport), COOLANG (PID2021-122263OB-C22), and CORTEX (PID2021-123956OB-I00) with funds granted by MCIN/AEI/10.13039/501100011033 and, as appropriate, by “ERDF A way of making Europe”, by the “European Union” or by the “European Union NextGenerationEU/PRTR”.
+on this paper was partially funded by the University of Alicante, the Spanish Ministry of Science and Innovation, the Generalitat Valenciana, and the European Regional Development Fund (ERDF) through the following funding: As main, VIVES: ‘Pla de Tecnologies de la Llengua per al valencià’ project (2022/TL22/00215334) from the Projecte Estratègic per a la Recuperació i Transformació Econòmica (PERTE) and the following supportive projects: NL4DISMIS (CIPROM/2021/21) with funds granted by the Generalitat Valenciana (Conselleria d’Educacio, Investigacio, Cultura i Esport), COOLANG (PID2021-122263OB-C22), and CORTEX (PID2021-123956OB-I00) with funds granted by MCIN/AEI/10.13039/501100011033 and, as appropriate, by ‘ERDF A way of making Europe’, by the ‘European Union’ or by the ‘European Union NextGenerationEU/PRTR’.
 Limitations
-While we believe that our study provides valuable insights into the measurement of gender bias in gendered languages using LLMs, we have identified several limitations.
+While we believe that our study provides valuable insights into the measurement of gender representation bias in gendered languages using LLMs, we have identified several limitations.
 Inclusion of epicene words: In our analysis, we included epicene words – nouns that have the same form, including grammatical gender, for male and for female referents.
@@
 This inherent ambiguity can lead to inaccuracies in gender classification.
-In the data analyzed in this work, epicene words represent 5.8 % of all identified person-referencing words.
+In the data analyzed in this work, epicene words represent only 5.8 % of all identified person-referencing words.
 It is noteworthy that 73.7 % of epicene words were counted as feminine (see Appendix C), such that excluding the epicene words would make the gender ratios even more biased towards the masculine gender.
-Contextuality of the analysis: Our methodology focuses primarily on the occurrence and classification of gendered nouns and pronouns without deeply analyzing the context in which these words appear.
+The role of context: Our methodology focuses primarily on the occurrence and classification of gendered nouns and pronouns without deeply analyzing the context in which these words appear.
 This approach provides a quantitative measure of gender representation but lacks the nuance of qualitative analysis.
 Understanding the context could reveal more about the nature of gender bias, such as whether certain roles or characteristics are predominantly associated with a specific gender.
-Future work should incorporate contextual analysis to provide a richer and more comprehensive understanding of gender bias in texts.
+Future work should incorporate contextual analysis to provide a richer and more comprehensive understanding of gender bias in text.
 Binary gender: The current framework represents gender as a binary variable – masculine and feminine – and does not account for non-binary or other gender identities.
@@
 As language evolves and becomes more inclusive, future studies should consider expanding the classification system to encompass a wider range of gender identities, ensuring that NLP models are sensitive to and representative of all individuals.
-Budget implications: Using commercial, state-of-the-art LLMs in research can have a non-negligible financial cost.
-While we selected representative subsets from widely used benchmark datasets, a more extensive analysis could potentially provide more robust and generalizable findings at a cost.
 LLM dependence and model variability: Our method heavily relies on the capabilities of advanced LLMs, specifically gpt-4-turbo.
@@
 Additionally, the use of proprietary models may impact the reproducibility and transparency of our findings.
+Budget considerations: Using commercial, state-of-the-art LLMs in research can have a non-negligible financial cost.
+While we selected representative subsets from widely used benchmark datasets, a more extensive analysis could potentially provide more robust and generalizable findings at a cost.
 Cultural and linguistic variations: Gender representation in language can be influenced by cultural and regional variations, which our study does not explicitly address.
@@
 By addressing these limitations in future work, we aim to refine our methodology, incorporate a broader range of linguistic and cultural nuances, and enhance the fairness and accuracy of gender bias detection in NLP systems.
-Ethics Statement
-We aim to promote fairness and inclusivity by identifying and quantifying biases in text corpora used to train LLMs.
-We have adhered to ethical standards and code of conduct by ensuring transparency, reproducibility, and validating our methodology against manually annotated data.
-The datasets used for evaluation are publicly available benchmark datasets.
-While our work highlights significant gender disparities, we recognize the limitations of focusing on grammar-based gender classification and the reliance on specific LLMs.
-We are committed to ethical AI use, advocating for continuous improvement in bias detection and mitigation techniques.
-Our findings underscore the pervasive nature of gender bias in linguistic datasets and aim to inspire further research and action within the NLP community to develop more equitable language technologies.
-By raising awareness and providing robust methodologies, we contribute to the broader discourse on AI ethics and the pursuit of technologies that uphold human dignity and equality.
 Appendix A Appendix: Prompt Formulation